OpenStudy (mecharv):

For the cardioids r = a(1+cosθ) , prove that (radius of curvature)^2/r is constant

OpenStudy (holsteremission):

Curvature for a polar curve defined by \(r(\theta)\) is \[\kappa (\theta) = \frac{\bigg|2 f'(\theta)^2 + f(\theta)^2 - f(\theta)f''(\theta)\bigg|}{\bigg[f'(\theta)^2 + f(\theta)^2 \bigg]^{3/2}}\]

OpenStudy (holsteremission):

Whoops, those \(f\)s should be \(r\)s. First compute your derivatives: \[r(\theta)=a(1+\cos\theta)\\[1ex] \quad\implies r'(\theta)=-a\sin\theta\\[1ex] \quad\implies r''(\theta)=-a\cos\theta\]

OpenStudy (mecharv):

I am aware of that I have also proceeded with the same... but im struggling to bring it to a constant value.

OpenStudy (mecharv):

That is the radius of curvature has a formula[\[\rho=\frac{ (1+y _{1}^{2})^{3/2} }{ y _{2^{2}}}\]

OpenStudy (holsteremission):

Ah okay. You have \[\begin{align*} \kappa(\theta)&=\frac{\bigg|2(-a\sin\theta)^2+(a(1+\cos\theta))^2-a(1+\cos\theta)(-a\cos\theta)\bigg|}{\bigg[(-a\sin\theta)^2+(a(1+\cos\theta))^2\bigg]^{3/2}}\\[1ex] &=\frac{\bigg|2a^2\sin^2\theta+a^2+2a^2\cos\theta+a^2\cos^2\theta+a^2\cos\theta+a^2\cos^2\theta\bigg|}{\bigg[a^2\sin^2\theta+a^2+2a^2\cos\theta+a^2\cos^2\theta\bigg]^{3/2}} \end{align*}\]Every instance of \(a^2\sin^2\theta+a^2\cos^2\theta\) reduces to \(a^2\), so you can simplify this to \[\begin{align*} \kappa(\theta)&=\frac{\bigg|3a^2+3a^2\cos\theta\bigg|}{\bigg(2a^2+2a^2\cos\theta\bigg)^{3/2}}\\[1ex] &=\frac{3a^2\bigg|1+\cos\theta\bigg|}{(2a^2)^{3/2}\bigg(1+\cos\theta\bigg)^{3/2}}\\[1ex] &=\frac{3}{\sqrt8|a|(1+\cos\theta)^{1/2}} \end{align*}\]The radius of curvature is given by \[\rho(\theta)=\frac{1}{\kappa(\theta)}\]You need to show that \(\dfrac{\rho(\theta)^2}{r}\) is constant. Can you continue from here?

OpenStudy (mecharv):

Yep, I got there too just now :) I'll complete the problem. Thanks a ton :)

OpenStudy (mecharv):

Man, how do you type so many equations so fast. I need to get a hang of that. I'll try to attach a pic :)

OpenStudy (holsteremission):

Lots of practice and a penchant for having things look nice :)

OpenStudy (mecharv):

\[\frac{ \rho(\theta)^2 }{ r}= \frac{ 8 a}{ 9 }\]

OpenStudy (holsteremission):

Looks good!

OpenStudy (mecharv):

Thanks a lot :)

OpenStudy (holsteremission):

You're welcome, happy to help